Find constants a and b in the function f(x)=axe^(bx)such that f(1/4)=1 and the function has a local maximum at x=1/4.
1
Expert's answer
2019-12-09T11:56:38-0500

Given, "f(x) = axe^{bx}"

"f(1\/4)=1"

"(a\/4)e^{b\/4}=1 ----(1)"

Local maximum at x=1/4, that means "f'(1\/4)=0"

"f'(x) = abxe^{bx} + ae^{bx}"


"f'(1\/4)= (ab\/4)e^{b\/4}+ae^{b\/4}=0 --(2)"

Solving the 2nd equation,

"e^{b\/4}(ab\/4+a)=0"

Now, "e^{b\/4}\u22600"

So, "(ab\/4 +a)=0"

Or, "a(b\/4+1)=0"

Now from equation (1), value of 'a' cannot be zero.

So, "(b\/4+1)=0"

Or, "b=-4"

Putting the value of 'b' in equation (1),

"(a\/4) e^{-1}=1"

"a\/4e=1"

Or, "a=4e"

So, "a=4e,b= -4"